40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 103 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  11911   Tue Jan 5 15:48:16 2016 SteveUpdatesafetyRat trap locations

Please look around when working close to these five locations. Use flashlights or leave lights on.

These mechanial traps are HAZARDOUS !

No visitors or tours till Monday, Jan 11  2016

Attachment 1: Control_Room.jpg
Control_Room.jpg
Attachment 2: EarmCESdoor.jpg
EarmCESdoor.jpg
Attachment 3: 1Y1.jpg
1Y1.jpg
Attachment 4: MC2.jpg
MC2.jpg
Attachment 5: XarmMaglev.jpg
XarmMaglev.jpg
  11910   Tue Jan 5 13:17:06 2016 ranaUpdateLSCAUX X Freq Noise attempt

The problem here is that the MC displacement noise is leading to large frequency excursions of the PSL beam. Options

  1. Feed back the low frequency PLL control signal to the MC2 length to suppress the excursion required by the Marconi. This is better than driving the laser, since the drive to the laser would be squashed by the MC locking loop.
  2. Put the beat signal through a divider? Don't know if this makes the Marconi more able to handle it.
  3. Turn on the MCL path. this will make the low frequency MC error signal go to the MC length, thereby reducing the low frequency feedback to the NPRO.
  11909   Tue Jan 5 09:48:52 2016 SteveUpdatePEMRat trap moved

Our janitor confirmed that Q was not hallucinating about this animal. The dropping size indicating a good size one in the IFO room.

One of the mechanical traps moved from the control room to the east arm, close to the " machine shop " door.

I'm going to get more traps.

Quote:

Two mechanical and two sticdky traps were set to catch univited visitor.

Absolutely no food or food remains into inside garbage cans!!!!!!!!!!!!!!!!!!!!!!!!!

Quote:

A small rat / large mouse just ran through the control room. Ugh.

 

 

  11908   Tue Jan 5 02:54:38 2016 ericqUpdateLSCAUX X Freq Noise attempt

[ericq, Gautam]

We set out to lock a marconi to the IR fiber beat of PSL + AUX X to measure some frequency noise, and failed.

In short, the Marconi's 1.6MHz max external FM isn't enough oomph to stabilize the PLL error signal. It's actually evident on the Agilent that the beat moves around a few times more than that, which I should've noticed sooner... We could briefly "lock" the PLL for a few tenths of a second, but weren't able to get a spectrum from this.

We also tried using the digital phase tracker temperature servo for some help at ~DC; this worked to the extent that we didn't have to twiddle the Marconi carrier frequency to stay on top of the fringes as the beat wandered, but it didn't otherwise stabilize the beat enough to make a difference in locking the PLL.

I suppose one more thing to try is to lock the PSL laser itself to each AUX laser in turn via PLL, and look for different / excess noise.

The Green and IR beat electronics are a in a little bit of disarray at the moment, but it's not like anyone else is going to be using them for the time being...

  11907   Mon Jan 4 16:45:11 2016 gautamUpdateGreen LockingY-end AUX PDH noise breakdown

Summary:

I've re-measured the noise breakdown for the Y-end AUX PDH system. Spectra are attached. I've also measured the OLTF of the PDH loop, from which the UGF appears to be ~8.5kHz. 

Discussion:

As Eric and Koji pointed out, the spectra uploaded here were clearly wrong as there were breaks in the spectra between decades of frequency. I redid the measurements, this time being extra careful about impedance mismatch effects. All measurements were made from the monitor points on the PDH box, which according to the schematic found here, have an output impedance of 49.9 ohms. So for all measurements made using the SR785 which has an input impedance of 1Mohm, or those which had an SR560 in the measurement chain (also high input impedance), I terminated the input with a 50ohm terminator so as to be able to directly match up spectra measured using the two different analyzers. I'm also using my more recent measurement of the actuator gain of the AUX laser to convert the control signal from V/rtHz to Hz/rtHz in the plotted spectra. 

As a further check, I locked the IR to the Y-arm by actuating on MC2, and took the spectrum of the Y-arm mirror motion using the C1CAL model. We expect this to match up well with the in-loop control signal at low frequencies. However, though the shapes seem consistent in Attachment #2 (light orange and brown curves), I seem to be off by a factor of 5- not sure why. In converting the Y-arm mirror motion spectrum from m/rtHz to Hz/rtHz, I multiplied the measured spectrum by \frac{3.907*10^6}{0.5*532*10^{-9}}, which I think is the correct conversion factor (FSR/(0.5*wavelength))?

Attachment 1: ErrSigBreakdown.pdf
ErrSigBreakdown.pdf
Attachment 2: controlSigBreakdown.pdf
controlSigBreakdown.pdf
Attachment 3: YEnd_PDH_OLTF.pdf
YEnd_PDH_OLTF.pdf
  11906   Mon Jan 4 16:09:54 2016 gautamUpdateGreen LockingY end laser (Lightwave) PZT calibration

Summary:

I redid this measurement and have now determined the actuator gain to be 4.61 +/- 0.10 MHz/V. This is now pretty consistent with the expected value of ~5MHz/V as reported here.

Details:

I made the following changes to the old methodology:

  1. Instead of integrating around the excitation frequency, I am now just taking the ratio of peak heights (phase tracker output / error signal monitor) to determine the actuator gain.
  2. I had wrongly assumed that the phase tracker output was calibrated to green Hz and not IR Hz, so I was dividing by two where this was not necessary. I think this explains why my previous measurement yielded an answer approximately half the expected value.

I also took spectra of the phase tracker output and error signal to make sure I was choosing my excitation frequencies in regions where there were no peaks already present (Attachment #1).

The scatter of measured actuator gains at various excitation frequencies is shown in Attachment #2.

Attachment 1: choosingExcFreqs.pdf
choosingExcFreqs.pdf
Attachment 2: laserPZTcalib.pdf
laserPZTcalib.pdf
  11905   Mon Jan 4 14:45:41 2016 rana, eq, kojiConfigurationComputer Scripts / Programsnodus pwd change

We changed the password for controls on nodus this afternoon. We also zeroed out the authorized_keys file and then added back in the couple that we want in there for automatic backups / detchar.

Also did the recommended Ubuntu updates on there. Everything seems to be going OK so far. We think nothing on the interferometer side cares about the nodus password.

We also decided to dis-allow personal laptops on the new Martian router (to be installed soon).

  11904   Wed Dec 30 11:09:52 2015 SteveUpdateSUS4.4M EQ

Suspensions recovered after 4.4 Mag EQ

 

Attachment 1: 4.4MeqDevore.png
4.4MeqDevore.png
Attachment 2: free.png
free.png
  11903   Mon Dec 28 14:12:23 2015 SteveUpdateSUS4.2M EQ

Suspensions are recovered after 4.2 Mag earth quake. No obvoius sign of damage.

 

Attachment 1: 4.2Meq.png
4.2Meq.png
Attachment 2: recovered4.2M.png
recovered4.2M.png
  11902   Sat Dec 26 10:34:43 2015 SteveUpdateGeneraltoday

PMC locked manually and PRM sus damping restored.

Attachment 1: 1d.png
1d.png
  11901   Wed Dec 23 16:15:47 2015 ranaOmnistructureALARMfire alarm

Fire alarm went off several minutes ago. Talked to security and they said there was no fire. It beeped twice again just now. No one has been working on the IFO today.

  11900   Wed Dec 23 15:43:02 2015 ranaSummaryPSLPMC FSS IMC RF summing box

The EOM upstream of the PMC is used as the phase corrector for the FSS/IMC servo. It is also used to apply the 35.5 MHz PDH RF sidebands for the PMC locking. There is a Pomona box which is used to merge the two signals onto a single cable for the EOM.

Does this circuit make sense to anyone?

Attachment 1: PMCFSSsummingBox.png
PMCFSSsummingBox.png
  11899   Wed Dec 23 03:27:04 2015 ranaUpdateComputer Scripts / ProgramsLHO EPICS slow down

https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=24321

This LHO log indicates that EPICS slow down could be due to NFS activity. Could we make some trend of NFS activity on Chiara and then see if it correlates with EPICS flatlines?

I wonder if our EPICS issues frequency is correlated to the Chiara install.

  11898   Tue Dec 22 16:44:03 2015 gautamUpdateGeneralFS725 Rubidium reference - REPAIRED
Quote:

However, the one that was recovered from the ATF lab is no longer locking to the Rubidium reference frequency, although it was locked at the time we disconnected it from the ATF lab. I emailed the support staff at SRS, who seem to think that either the internal oscillator has drifted too far, or the Rb lamp is dead. Either ways, it needs to be repaired. They suggested that I run a check by issuing some serial commands to the unit to determine which of these is actually the problem, but I've been having some trouble setting up the serial link - I will try this again tomorrow.

The Rubidium standard we had sent in for repair and recalibration has come back. I checked the following:

  • Powered the unit on - it was locked to the internal rubidium reference within a few minutes as prescribed in the manual.
  • After it had locked to the internal reference, I checked that it was able to lock to an external 1pps reference from our GPS timing unit- this too was achieved within a few minutes as prescribed in the manualyes

However, I am still having trouble setting up a serial communications link with the FS725 with a USB-serial adaptor - I've tried with a Raspberry Pi and my Mac (using screen to try and connect), and also using one of the old Windows laptops lying around on which I was able to install the native software supplied by SRS (still using the USB-serial adaptor to establish connection though). Could it be that the unit is incompatible with the USB-serial adaptor? I had specifically indicated in the repair request that this was also a problem. In any case, this doesn't seem to be crucial, though it would have been nice for diagnostics purposes in the future...

I've stored the repaired FS725 inside the electronics cabinet (marked "Eletronics Modules") for now (the other unit was returned to Antonio in W. Bridge some weeks ago). 

Attachment 1: FS725_repaired.jpg
FS725_repaired.jpg
  11897   Tue Dec 22 16:39:42 2015 SteveUpdateendtable upgradeETMY optical table enclosure

I think there should be a scientifically based aveluation of the ETMY enclosure so we can make the ETMX better.

Meanwhile I'm counting pieces to move on with the south end table cover.

 

Attachment 1: IMG_0028.JPG
IMG_0028.JPG
Attachment 2: IMG_0027.JPG
IMG_0027.JPG
Attachment 3: ETMY-ISCT_EISOL.jpg
ETMY-ISCT_EISOL.jpg
  11896   Tue Dec 22 16:23:33 2015 gautamUpdateIOOInput alignment to PMC tweaked

When I came in this afternoon, I saw that the PZT voltage to the PMC had railed. Following the usual procedure of turning the servo gain to zero and adjusting the DC offset, I got the PMC to relock, but the PMCR level was high and the alignment looked poor on the control room monitor. So I tweaked the input alignment on the PSL till I felt it was more reasonable. The view on the control room monitor now looks more like the usual state, and the "REFL (V)" field on the PMC MEDM screen now reads 0.02-0.03 which is the range I remember it being in nominally. 

  11895   Mon Dec 21 14:31:41 2015 SteveUpdatePEMRat traps set

Two mechanical and two sticdky traps were set to catch univited visitor.

Absolutely no food or food remains into inside garbage cans!!!!!!!!!!!!!!!!!!!!!!!!!

Quote:

A small rat / large mouse just ran through the control room. Ugh.

 

Attachment 1: ratsNC.jpg
ratsNC.jpg
  11894   Mon Dec 21 02:29:49 2015 ericqUpdateLSCAUX X RIN measurements

I'll finish up the beat / frequency noise parts of the diagnosis tomorrow later, but I've done some investigation of the AUX X laser RIN. 

I placed a PDA255 at one of the rejected beams from the PBS on the downstream side of the IR faraday, making sure the power didn't saturate the PD. I measured the RIN on a SR785, and simultaneously looked at the signal on a 100MHz scope. 

The RIN has a very strong dependence on the laser diode current, and no noticable dependence on the crystal temperature or the presence of the PDH modulation / temperature control cables. Here are some traces, note that "nominal" current up until recently was 2.0A. 

When adjusting the diode current, a peak beings to appear in the tens of kHz, eventually noticible in the DC power trace on the scope. The point at which this occurs is not fixed.

At all times, I saw a strong intensity fluctuation at around 380-400kHz on the scope whose amplitude fluctuated a fair amount (at least 75mVrms over Vdc=6.5V, but would often be 2 or 3 times that).

I didn't look at the frequency noise while doing this, because the WiFi at the X end was too slow, I'll do more tomorrow in the daytime. 

Attachment 1: auxXRIN.pdf
auxXRIN.pdf
  11893   Sun Dec 20 23:23:54 2015 ericqUpdateALARMRats.

A small rat / large mouse just ran through the control room. Ugh.

  11892   Fri Dec 18 17:37:04 2015 ranaUpdateLSCUncooperative AUX X

Here's how we should diagnose the EX laser:

  1. Compare IR RIN of laser out to 100 kHz with that of another similar NPRO.
  2. Look at time series of IR beat signal with a fast scope. Are there any high frequency glitches?
  3. Disconnect all of the cables to the EX laser PZT and temperature control. Does the frequency noise change?
  4. Change the temperature by +/- 1 deg to move away from mode hop regions. Remeasure RIN and frequency noise and plot.
  11891   Thu Dec 17 16:44:03 2015 gautamUpdateCDSALS Slow control MEDM screen updated
Quote:

I've not updated the MEDM screens to reflect the two new paths yet, but will do so soon. It also remains to install appropriate filters for the servo path that takes the frequency readout as the input.

A few more related changes:

  1. The couplers that used to sit on the green beat PDs on the PSL table have now been shifted to the IR broadband PDs in the FOL box so that I can get the IR beat frequency over to the frequency counters. The FOL box itself, along with the fibers that bring IR light to the PSL table, have been relocated to the corner of the PSL table where the green beat PDs sit because of cable length constraints.
  2. I've updated the ALS slow control MEDM screen to allow for slow control of the beat frequency. The servo shape for now is essentially just an integrator with a zero at 1 Hz. The idea is to set an offset in the new filter module, which is the desired beat frequency, and let the integrator maintain this beat frequency. One thing I've not taken care of yet is automatically turning this loop off when the IMC loses lock. Screenshot of the modified MEDM screen is attached. 
  3. I checked the performance by using the temperature sliders to introduce an offset. The integrator is able to bring the beat frequency back to the setpoint in a few seconds, provided the step I introduced was not two big (~20 counts, but this is a pretty large shift in beat frequency, nearly 20MHz).

To do:

  1. Figure out how to deal with the IMC losing lock. I guess this is important if we want to use the IR beatnote as a diagnostic for the state of the X AUX laser.
  2. Optimize the servo gains a little - I still see some ringing when I introduce an offset, this could be avoided...
Attachment 1: ALS_SLOW_17DEC2015.png
ALS_SLOW_17DEC2015.png
  11890   Thu Dec 17 14:02:05 2015 gautamUpdateCDSIPC channels for beat frequency control set up

I've set up two IPC channels that take the output from the digital frequency counters and send them to the end front-ends (via the RFM model). A summary of the steps I followed:

  1. Set up two Dolphin channels in C1ALS to send the output of the frequency counter blocks to C1RFM (I initially used RFM blocks for these, but eric suggested using Dolphin IPC for the ALS->RFM branch, as they're faster.. Eric's removed the redundant channel names)
  2. Set up two RFM channels in C1RFM to send the out put of the frequency counter blocks to C1SCX/Y (along with CDS monitor points to monitor the error rate and a filter module between the ALS->RFM and RFM->SCX/Y IPC blocks - I just followed what seemed to be the convention in the RFM model).
  3. Set up the receiving channels in C1SCX and C1SCY
  4. Re-compiled and re-started the models in the order C1ALS, C1RFM, C1SCX and C1SCY.

I've set things up such that we can select either the "PZT IN" or the frequency counter as the input to the slow servo, via means of a EPICS variable called "FC_SWITCH" (so C1:ALS-X_FC_SWITCH or C1:ALS-Y_FC_SWITCH). If this is 0, we use the default "PZT IN" signal, while setting it to 1 will change the input to the slow servo to be the frequency readout from the digital frequency counter. I've not updated the MEDM screens to reflect the two new paths yet, but will do so soon. It also remains to install appropriate filters for the servo path that takes the frequency readout as the input.

Tangentially related to this work: I've modified the FC library block so that it outputs frequency in MHz as opposed to Hz, just for convenience..

  11889   Thu Dec 17 01:55:16 2015 ericqUpdateLSCUncooperative AUX X

[ericq, Gautam]

We were not able to fix the excess frequency noise of the AUX X laser by the usual laser diode current song and dance. Unfortunately, this level of noise is much too high to have any realistic chance of locking.  angry

We're leaving things back in the IR beat -> phase tracker state with free running AUX lasers, on the off chance that there may be anything interesting to see in the overnight data. This may be limited by our lack of automatic beatnote frequency control. (Gautam will soon implement this via digital frequency counter). I've upped the FINE_PHASE_OUT_HZ_DQ frame rate to 16k from 2k, so we can see more of the spectrum.

For the Y beat, there is the additional weird phenomenon that the beat amplitude slowly oscillates to zero over ~10 minutes, and then back up to its maximum. This makes it hard for the phase tracker servo to stay stable... I don't have a good explanation for this. 

  11888   Wed Dec 16 23:15:28 2015 ericqUpdateGreen LockingGreen beat channels temporarily set up as IR beat channels

With the IR beats going to the nominal ALS channels as Gautam left them, we're able to measure the free running frequency noise of the end AUX lasers. 

Specifically, the end shutters are closed, leaving the AUX lasers free running. The IR beats then consist of this free running light beating with the PSL light, and the ALS phase trackers give a calibrated frequency noise spectrum. I've stabilized the PSL light by locking the laser to the Y arm via MC2 acutation, so the free running AUX laser noise should dominate by a lot above the suspension resonances. This also has the benefit of giving me the use of the CAL'd Y arm displacement as a sanity check. 

At this point in time, it looks like the X laser is close to 10x noisier than the Y laser, though it does seem to be at the rule-of-thumb "10kHz/rtHz at 100Hz" level. 

Attachment 1: 2015-12-16_AUXfreerunning.pdf
2015-12-16_AUXfreerunning.pdf
  11887   Wed Dec 16 18:34:40 2015 gautamUpdateGreen LockingGreen beat channels temporarily set up as IR beat channels

Since there are a few hours to go before the locking efforts tonight, I've temporarily borrowed the channels used to read out the green beat frequency, and have hooked them up to the broadband IR PDs in the FOL box on the PSL table. I've used the network analyzer in the control room to roughly position the two beatnotes. I've also turned the green beat PDs back on (since the PSL shutter has to be open for the IR beat, and there is some green light falling on these PDs, but I've terminated the outputs).

So this needs to be switched back before locking efforts tonight...

  11886   Wed Dec 16 10:56:22 2015 gautamUpdateCDShard reboot of FB

[ericq,gautam]

Forgot to submit this yesterday...

While we were trying to get the X-arm locked to IR using MC2, frame-builder mysteriously crashed, necessitating us having to go down to the computer and perform a hard reboot (after having closed the PSL shutter and turning all the watchdogs to "shutdown"). All the models restarted by themselves, and everything seems back to normal now..

  11885   Wed Dec 16 10:22:14 2015 SteveUpdateIOOthis morning

c1sus and c1ioo were restarted. PMC locked.
 

Attachment 1: PMClocked.png
PMClocked.png
  11884   Tue Dec 15 18:08:22 2015 Max IsiUpdateGeneralSummary archive cleaning cron job

I have added a new cron job in pcdev1 at CIT using the 40m shared account. This will run the /home/40m/DetectorChar/bin/cleanarchive script one minute past midnight on the first of every month. The script removes GWsumm archive files older than 1 month old.

  11883   Tue Dec 15 11:22:53 2015 gautamUpdateCDSc1scx and c1asx crashed

I noticed what I thought was excessive movement of the beam spot on ITMX and ETMX on the control room monitors, and when I checked the CDS FE status overview MEDM screen, I saw that c1scx and c1asx had crashed. I ssh-ed into c1iscex and restarted both models, and then restarted fb as well. However, the DAQ-DCO_C1SCX_STATUS indicator remains red even after restarting fb (see attached screenshot). I am not sure how to fix this so I am leaving it as is for now, and the X arm looks to have settled down.

Attachment 1: CDS_FE_STATUS_OVERVIEW_15DEC2015.png
CDS_FE_STATUS_OVERVIEW_15DEC2015.png
  11882   Mon Dec 14 23:56:29 2015 ericqUpdateCDSc1pem reverted

To get C1PEM data back into the frames, I removed the new BLRMS blocks, recompiled, reinstalled, re-enabled it in daqd, restarted.

We still really want more headroom in our framebuilder situation. 

  11881   Mon Dec 14 23:49:03 2015 ericqUpdateOptical LeversCalibration of oplevs for ITMX/ETMX
Quote:

Based on calibration measurement I have done (elog 11785, 11831), I updated calibration factors of oplevs on medm screen as follows. Not to change loop gain oplev servo, I also changed oplev servo gain.

After making sure that the upper UGFs were properly in place, I saved these settings to the SDF files. Thanks Yutaro!

  11880   Mon Dec 14 16:46:42 2015 ericqUpdateWienerFilteringNoise Subtraction Puzzler

Here's something to ponder.

Our online MCL feedforward uses perpendicular vertex T240 seismometer signals as input. When designing a feedforward filter, whether FIR Wiener or otherwise, we posit that the PSD of the best linear subtraction one can theoretically achieve is given by the coherence, via Psub = P(1-C). 

If we have more than one witness input, but they are completely uncorrelated, then this extends to Psub = P(1-C1)(1-C2). However, in reality, there are correlations between the witnesses, which would make this an overestimate of how much noise power can be subtracted. 

Now, I present the actual MCL situation. [According to Ignacio's ELOG (11584), the online performance is not far from this offline prediction]

Somehow, we are able to subtract much more noise at ~1Hz than the coherence would lead you to believe. One suspicion of mine is that the noise at 1Hz is quite nonstationary. Using median [C/P]SDs should help with this in principle, but the above was all done with medians, and using the mean is not much different.

Thinking back to one of the metrics that Eve and Koji were talking about this summer, (std(S)/mean(S), where S is the spectrogram of the signal) gives an answer of ~2.3 at that peak at 1.4Hz, which is definitely in the nonstationary regieme, but I don't have much intution into just how severe that value is.

So, what's the point of all this? We generally use coherence as a heuristic to judge whether we should bother attempting any noise subtraction in the first place, so I'm troubled by a circumstance in which there is much more subtraction to be had than coherence leads us to believe. I would like to come up with a way of predicting MISO subtraction results of nonstationary couplings more reliably.

Attachment 1: subpuzz.pdf
subpuzz.pdf
  11879   Mon Dec 14 16:27:11 2015 gautamUpdateGreen LockingY-end AUX PDH noise breakdown

Summary:

I've attached the results from my measurements of the noise characteristics of the Y-end auxiliary PDH system.

Details:

The following spectra were measured, in the range DC-1MHz:

  1. Analyzer noise floor (measured with input terminated)
  2. Green REFL PD dark noise (measured with the Y-end green shutter closed)
  3. Mixer noise (measured with input to mixer terminated - measured with an SR560 with a gain of 100)
  4. Servo noise (measured with input to servo terminated)
  5. In loop error signal (measured with green locked to Y-arm, LSC off - using monitor point on PDH box)
  6. In loop control signal (measured with green locked to Y-arm, LSC off - using monitor point on PDH box)

In order to have good spectral resolution, the frequency range was divided into 5 subsections: DC-200Hz, 200Hz-3.4kHz, 3.4kHz-16.2kHz, 10kHz-100kHz, 100kHz-1MHz. The first three are measured using the SR785, while the last two ranges are measured with the Agilent network analyzer. The spectrum of the mixer output with its input terminated was quite close to the analyzer noise floor - hence, this was measured with an S560 preamplifier set to a gain of 100, and subsequently dividing the ASD by 100. To convert the Y-axis from V/rtHz to Hz/rtHz, I used two conversion factors: for the analyzer noise floor, PD dark noise, mixer noise and in-loop error signal, I made an Optickle simulation of a simple FP cavity (all parameters taken from the wiki optics page, except that I put in Yutaro's measured values for the arm loss and a modulation depth of 0.21 which I estimated as detailed here), and played around with the demodulation phase until I got an error signal that had the same qualitative shape as what I observed on an oscilloscope with the arms freely swinging (feedback to the laser PZT disabled). The number I finally used is 45.648 kHz/V (the main horns were 800mV peak-to-peak on an oscilloscope trace, results of the Optickle FP cavity simulation shown in Attachment #2 used to calibrate the X-axis). For the servo noise spectrum and in-loop control signal, I used the value of 2.43 MHz/V as determined here

I'm not sure what to make of the strong peaks in the mixer noise spectrum between ~60Hz and 10kHz - some of the more prominent peaks are 60Hz harmonics, but there are several peaks in between as well (these have been confusing me for some time now, they were present even when I made the measurement in this frequency range using the Agilent network analyzer. My plan is to repeat these measurements for the Xend now. 

Attachment 1: YAUX_NB_Dec2015.pdf
YAUX_NB_Dec2015.pdf
Attachment 2: PDH_errSig_Calib.pdf
PDH_errSig_Calib.pdf
  11878   Mon Dec 14 14:08:49 2015 SteveUpdateSUSRuby wire standoffs update

Two companies are willing to  make the ruby grooves and the third one is still working on their quote.

The price is ~$100 each. The cost goes down 10% if we  order 50 instead of 30 pieces.

How many should we get ?

  11877   Sun Dec 13 21:55:28 2015 gautamUpdateGreen LockingY end laser (Lightwave) PZT calibration

Summary:

After the discussions at the Wednesday meeting, I redid this measurement using a sinusoidal excitation summed at the error-point of the PDH servo as opposed to a DC offset. From the data I collected, I measured the actuator gain to be 2.43 +/- 0.04 MHz/V. This is almost half the value we expect, I'm not sure if I'm missing something obvious.


Details:

  1. Attachment #1 is a sketch of the measurement setup and points at which signals are measured/calculated. Some important changes:
    • I am now using the channel C1:ALS-Y_ERR_MON_OUT to directly measure the input signal to the servo. In order to get the calibration constant for this channel from counts to volts, I simply hooked up the input to the channel to an oscilloscope and noted the amplitude of the signal seen on the scope in volts. The number I have used is 52uV/count (note that the signal to the ADC is amplified by a factor of 10 by an SR560).
    • I measured the transfer function from the input to the servo (marked "A" in the sketch) to the output of the Pomona box going to the laser PZT (marked "B" on the sketch) using an SR785 - see Attachment #2. This allowed me to convert the amplitude of excitation at A to an amplitude at B, which is what we need, as we want to measure C/B.
  2. The measurement itself was done by locking the arms to IR, running ASS to maximize IR transmission, setting up a green beat note, and then measuring the two channels of interest with the excitation to the error-point on. 
  3. I was initially trying to use time-series plots to measure these amplitudes - Koji suggested I use the Fourier domain instead, and so I took FFTs of the two channels we are interested in (using a flat-top window with 0.1 Hz BW) and estimated the RMS values at the frequency at which I had injected an excitation. Data+code used is in Attachment #3. In particular, I was integrating the PSD over 1Hz centered at the excitation frequency in order to calculate the RMS power at the excitation frequency - it could be that for C1:ALS-BEATY_FINE_PHASE_OUT_HZ, the spectral leakage into neighbouring bins is more significant than for C1:ALS-Y_ERR_MON_OUT (see Attachment #4)?
  4. With the amplitudes thus obtained, I took the ratio C/B (see sketch) to determine the MHz/V actuator gain. I had injected excitations at 5 frequencies (916Hz, 933Hz, 977Hz, 1030Hz and 1067Hz, choses in relatively "quiet" parts of the spectrum of C1:ALS-Y_ERR_MON_OUT with no excitations), and the result reported is the average from these five measurements, while the error is the standard deviation in the 5 measurements.
  5. Unrelated to this meaurement - while I had the SR560 hooked up to the input of the PDH box, I inverted the mixer output to the servo input, as I thought I could use this method to estimate the modulation depth. I did so by locking the Y arm green to the sideband TEM00 mode, and comparing the green transmission in this state to that when the Y arm is locked to a carrier TEM00 mode. I averaged C1:ALS-TRY_OUT for 10 seconds in 3 cases: (i) Carrier TEM00, (ii)sideband TEM00, and (iii) shutter closed - from this measurement, I estimate the modulation depth to be 0.209 +/- 0.006 (errors used to calculate the total error were the standard deviations of the measured transmission). 

Next steps:

  1. Check that I have not missed out anything obvious in estimating the actuator gain - particularly the spectral leakage bit I mentioned above.
  2. If this methodology and measurement is legitimate, repeat for the X end, and complete the noise budgeting for both AUX PDH loops.
Attachment 1: IMG_5972.JPG
IMG_5972.JPG
Attachment 2: ServoY_TF_13Dec2015.pdf
ServoY_TF_13Dec2015.pdf
Attachment 3: DatanCode.zip
Attachment 4: PSD_916Hz.pdf
PSD_916Hz.pdf
  11876   Fri Dec 11 23:12:09 2015 KojiSummaryCOCLoss map measurement document

Yutaro left detailed slides for his loss map measurement

https://dcc.ligo.org/LIGO-G1501547

  11875   Fri Dec 11 16:16:36 2015 yutaroUpdateOptical LeversCalibration of oplevs for ITMX/ETMX

Based on calibration measurement I have done (elog 11785, 11831), I updated calibration factors of oplevs on medm screen as follows. Not to change loop gain oplev servo, I also changed oplev servo gain.

C1:SUS-ETMX_OL_PIT_CALIB, C1:SUS-ETMX_OL_PIT_GAIN

(45.1,16) => (200,3.5)

C1:SUS-ETMX_OL_YAW_CALIB, C1:SUS-ETMX_OL_YAW_GAIN

(85.6,8) => (222,3.0) 

C1:SUS-ETMY_OL_PIT_CALIB, C1:SUS-ETMY_OL_PIT_GAIN

(26,-16) => (140,-3.0) 

C1:SUS-ETMY_OL_YAW_CALIB, C1:SUS-ETMY_OL_YAW_GAIN

(31,-21) => (143,-4.5) 

C1:SUS-ITMX_OL_PIT_CALIB, C1:SUS-ITMX_OL_PIT_GAIN

(110,8) => (122,7.2) 

C1:SUS-ITMX_OL_YAW_CALIB, C1:SUS-ITMX_OL_YAW_GAIN

(81,-11) => (147,-6) 

C1:SUS-ITMY_OL_PIT_CALIB, C1:SUS-ITMY_OL_PIT_GAIN

(159,15) => (239,10) 

C1:SUS-ITMY_OL_YAW_CALIB, C1:SUS-ITMY_OL_YAW_GAIN

(174,-21) => (226,-16) 

 

  11874   Fri Dec 11 15:37:50 2015 yutaroUpdateLSCPower recycling gain estimation from arm loss measurement

Attached is the plot of relation between the average arm round trip loss and power recycling gain. 2 % loss due to PR3 AR reflection is taken into account.

Attachment 1: PRG_plot.png
PRG_plot.png
  11873   Fri Dec 11 13:28:36 2015 KojiUpdateLSCPower recycling gain estimation from arm loss measurement

Can I ask you to make a plot of the power recycling gain as a function of the average arm loss, indicating the current loss value?

  11872   Fri Dec 11 09:35:44 2015 yutaroUpdateLSCPower recycling gain estimation from arm loss measurement

I took PR3 AR reflectivity and calculated PRG (PR3 is flipped and so AR surface is inside PRC).

As shown in attached figure, which shows AR specification of the LaserOptik mirror (PR3 is this mirror), AR reflectivity of PR3 is ~0.5 %. Since resonant light in PRC goes through AR surface of PR3 4 times per round trip, round trip loss due to this is ~2 %. Then I got

PRG = 7.8.    

 

Attachment 1: LaserOptikAR.png
LaserOptikAR.png
  11871   Thu Dec 10 19:53:22 2015 yutaroUpdateLSCstrange behavior of ASDC

To check if the strange behavior of ASDC is caused by SR2/SR3 or not, I did the following measurement:

ASDC measures the power of the light reflected by ITMX. POXDC measures the power of the light reflected by ITMX and SRM successively. Then I varied the angle of ITMX in YAW direction and compared the behaviors of ASDC and POXDC.

The results are shown in Attachments 1-3.

As you can see in these figures, the strange up-and-down behavior appeared ONLY in ASDC. Therefore, the cause of this behavior exists between AS table and SRM (I had confirmed that the angle of SRM did not affect ASDC).

And this behavior is fringe-like, as can be seen in the figures (there seems to be 3 "peaks" and 2 "valleys"), so the cause could be interference between main path and not good AR reflection at a mirror after SRM before AS table (I suspect a mirror is flipped mistakenly).   

Attachment 1: 30.png
30.png
Attachment 2: 11.png
11.png
Attachment 3: 49.png
49.png
  11870   Thu Dec 10 12:33:04 2015 yutaroUpdateLSCstrange behavior of ASDC

I did additional tests for the strange behavior of ASCD. ETMY, ETMX and ITMY were misaligned so that only light reflected by ITMX went into AS port. I had done similar measurement before with ITMY YAW varied.

Attachment 1 shows how ASDC level changed when ITMX PIT varied.

Attachment 2 shows how ASDC level changed when ITMX YAW varied.

Attachment 3 shows how the power of light measured by a power meter just after the AS view port varied when ITMX YAW varied.

 

Comparing 1 & 2, we can say that this behavior is not unique to YAW direction.

From 2 & 3, we can say something strange is happening inside the chamber.   

 

Attachment 1: 07.png
07.png
Attachment 2: 28.png
28.png
Attachment 3: ASDC.png
ASDC.png
  11869   Wed Dec 9 23:16:13 2015 ranaUpdateComputer Scripts / ProgramsNodus security

NDS2 and the usual ports so that we can use optimus as a comsol server.

Quote:

 

I don't think there are any other ports we need open, but I could be wrong. Let me know if I broke something you need!

 

  11868   Wed Dec 9 19:01:45 2015 jamieUpdateCDSback to fb1

I spent this afternoon trying to debug fb1, with very little to show for it.  We're back to running from fb.

The first thing I did was to recompile EPICS from source, so that all the libraries needed by daqd were compiled for the system at hand.  I compiled epics-3.14-12-2_long from source, and installed it at /opt/rtapps/epics on local disk, not on the /opt/rtapps network mount.  I then recompiled daqd against that, and the framecpp, gds, etc from the LSCSoft packages.  So everything has been compiled for this version of the OS.  The compilation goes smoothly.

There are two things that I see while running this new daqd on fb1:

instability with mx_streams

The mx stream connection between the front ends and the daqd is flaky.  Everything will run fine for a while, the spontaneously one or all of the mx_stream processes on the front ends will die.  It appears more likely that all mx_stream processes will die at the same time.  It's unclear if this is some sort of chain reaction thing, or if something in daqd or in the network itself is causing them all to die at the same time.  It is independent of whether or not we're using multiple mx "end points" (i.e. a different one for each front end and separate receiver threads in the daqd) or just a single one (all front ends connecting to a single mx receiver thread in daqd).

Frequently daqd will recover from this.  The monit processes on the front ends restart the mx_stream processes and all will be recovered.  However occaissionally, possibly if the mx_streams do not recover fast enough (which seems to be related to how frequently the receiver threads in daqd can clear themselves), daqd will start to choke and will start spitting out the "empty blocks" messages that are harbirnger of doom:

Aborted 2 send requests due to remote peer 00:30:48:be:11:5d (c1iscex:0) disconnected
00:30:48:d6:11:17 (c1iscey:0) disconnected
mx_wait failed in rcvr eid=005, reqn=182; wait did not complete; status code is Remote endpoint is closed
disconnected from the sender on endpoint 005
mx_wait failed in rcvr eid=001, reqn=24; wait did not complete; status code is Remote endpoint is closed
disconnected from the sender on endpoint 001
[Wed Dec  9 18:40:14 2015] main profiler warning: 1 empty blocks in the buffer
[Wed Dec  9 18:40:15 2015] main profiler warning: 0 empty blocks in the buffer
[Wed Dec  9 18:40:16 2015] main profiler warning: 0 empty blocks in the buffer

My suspicion is that this time of failure is tied to the mx stream failures, so we should be looking at the mx connections and network to solve this problem.

frame writing troubles

There's possibly a separate issue associated with writing the second or minute trend files to disk.  With fair regularity daqd will die soon after it starts to write out the trend frames, producing the similar "empty blocks" messages.

  11867   Wed Dec 9 18:45:58 2015 KojiSummaryGeneralNetwork Topology Check

[Eric Q, Gautam, Koji]

We went through the network connections to produce the mapping of the instruments.
Gautam summarized the notes into a spread sheet. See attachments.

We didn't find any irregular connections except for the connection of NETMGR port of c1ioo to Martian Network.
This cable was removed.

Attachment 1: Network_topology_9Dec2015.xlsx
Attachment 2: Network_topology_9Dec2015.pdf
Network_topology_9Dec2015.pdf Network_topology_9Dec2015.pdf Network_topology_9Dec2015.pdf
  11866   Wed Dec 9 11:25:55 2015 SteveUpdateVACnormal RGA scan at day 435-436

Glitches are gone. Rga scan is good

 

Attachment 1: glichesGone.png
glichesGone.png
Attachment 2: d436.png
d436.png
  11865   Tue Dec 8 23:24:08 2015 gautamUpdateGreen LockingY end laser (Lightwave) PZT calibration

Summary:

I measured the PZT actuator gain for the Lightwave NPRO at the Y-end to be 3.6 +/- 0.3 MHz/V. This is somewhat lower than the value of 5 MHz/V reported here, but I think is consistent with that measurement. 

Details:

In order to calibrate the Y-axis of my Aux PDH loop noise budget plots, I wanted a measurement of the end laser actuator gain. I proceeded to measure this as follows:

  1. Use a function generator to add a DC offset to the error point - I did this by taking the output of the RF mixer -> Input A of an SR560, output of the function generator -> input B of the SR560 (via a 20 Ohm attenuator, and with a 50ohm T-eed to the input for impedance matching), and setting the output to A-B, and feeding that to the "Servo Input" on the PDH box.
  2. I then locked the arm to IR, ran the dither to maximize the green transmission, and set up a beat note at ~39 MHz with the help of the analyzer in the control room.
  3. Set phase tracker UGF, clear phase history.
  4. Vary the DC offset to the error point by using the offset on the function generator. I varied the offset until the green TEM00 lock was lost, in steps of 0.1 V. At each step, I averaged the output of the phase tracker for 15 seconds.
  5. Convert the applied DC offset to the DC offset appearing at the servo output using the transfer function of the servo box (DC gain measured to be ~65 dB), taking into account the 20dB attenuator also.

The attached plot shows the measured data. The X-axis is shown after the conversion mentioned in the last bullet point. The error bars are the standard deviations of the averaging at each DC offset. 


To do:

  1. The value of the DC gain of the servo, 65 dB, is an approximate one based on a rough measurement I did earlier today. I'll take a TF measurement with an SR785 tomorrow, but I think this shouldn't change the number too much.
  2. Upload the noise budget measurements for the Y-end PDH loop.
Attachment 1: Ycalib_8Dec.pdf
Ycalib_8Dec.pdf
  11864   Tue Dec 8 15:57:16 2015 yutaroSummaryLSCPower recycling gain estimation from arm loss measurement

I estimated power recycling gain with the results of arm loss measurement.

From elog 11818 and 11857, round trip losses including transmittivity of ETM of Y arm and X arm (let us call them T_\mathrm{loss,Y} and T_\mathrm{loss,X}) are 229+13.7=243 ppm and 483+13.7=495 ppm, respectively.

 

How I calculated:

I used the following formula.

Amplitude reflectivity of an arm cavity r_\mathrm{FP}

r_\mathrm{FP}=\sqrt{1-\frac{4T_\mathrm{ITM}T_\mathrm{loss}}{T^2_\mathrm{tot}}}   (see elog 11816)

Amplitude reflectivity of FPMI r_\mathrm{FPMI}

r_\mathrm{FPMI}=\frac{1}{2}(r_\mathrm{FP,X}+r_\mathrm{FP,Y})

With power transmittivity of PRM T_\mathrm{PRM} and amplitude reflectivity of PRM r_\mathrm{PRM}, power recycling gain is

\mathrm{PRG}=\frac{T_\mathrm{PRM}}{(1-r_\mathrm{PRM}r_\mathrm{FPMI})^2}.

 I assumed T_\mathrm{ITM}\simeq T_\mathrm{tot}=\frac{2\pi}{401}=0.01566T_\mathrm{PRM}=0.05637, and r_\mathrm{PRM}=\sqrt{1-T_\mathrm{PRM}}, and then I got

PRG = 9.8.

Since both round trip losses have relative error of ~ 4 % and PRG is proportional to inverse square of T_\mathrm{loss} up to the leading order of it, relative error of PRG can be estimated as ~ 8 %, so PRG = 9.8 +/- 0.8

 

Discussion

According to elog 11691, which says TRX and TRY level was ~125 when DRFPMI was locked, power recycling gain was \mathrm{PRG}=125\times T_\mathrm{PRM}=7.0 at the last DRFPMI lock.

Measured PRG is lower than PRG estimated here, but it is natural because various causes such as mode mismatch between PRC mode and arm cavity mode, imperfect contrast of FPMI, and so on could decrease PRG, which Eric suggested to me. 

 

Added on Dec 9

If T_\mathrm{loss,X} were as small as T_\mathrm{loss,Y}, PRG would be 16.0. PRC would be still under coupled.  

  11863   Tue Dec 8 15:40:48 2015 SteveUpdateVACglitchy RGA scan at day 434

The noise floor of the Rga scan is glitching less today

 

Attachment 1: lessGlichingToday.png
lessGlichingToday.png
  11862   Tue Dec 8 15:18:29 2015 ericqUpdateComputer Scripts / ProgramsNodus security

I've done a couple things to try and make nodus a little more secure. Some have worried that nodus may be susceptible to being drafted into a botnet, slowing down our operations. 

1. I configured the ssh server settings to disallow logins as root. Ubuntu doesn't enable the root account by default anyways, but it doesn't hurt.

2. I installed fail2ban. Function: If some IP address fails to authenticate an ssh connection 3 times, it is banned from trying to connect for 10 minutes. This is mostly for thwarting mass brute force attacks. Looking at /var/log/auth.log doesn't indicate any of this kind of thing going on in the past week, at least.

3. I set up and enabled ufw (uncomplicated firewall) to only allow incoming traffic for:

  • ssh
  • ELOG
  • Nodus apache stuff (svn, wikis, etc.)

I don't think there are any other ports we need open, but I could be wrong. Let me know if I broke something you need!

ELOG V3.1.3-